Eight hours after a runner and cyclist leave the house at the same time and head in the same direction, the cyclist is 88 miles ahead of the runner. If the cyclist's speed is 2 mph more than twice the speed of runner, find the speed of the runner and the cyclist.
Found 2 solutions by jorel1380, ikleyn:
Answer by jorel1380(3719)

Let n be the speed of the runner. Then the cyclist's speed is 2n+2. So:
8(2n+2)-8n=88
16n+16-8n=88
8n=72
n=9
2n+2=20
☺☺☺☺

Answer by ikleyn(52787)

.
See the lesson
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

You will find there many other solved similar problems/samples of this type.